# DITING: A Static Analyzer for Identifying Bad Partitioning Issues in TEE Applications

### v1

@@ 2.2.3. Direct Usage of Shared Memory
Shared memory, as a zero-copy memory block, provides the same physical memory space accessible to both the secure and non-secure partitions. Since this memory is maintained by the normal-world code, bad partitioning can introduce significant vulnerabilities. If the TEE code directly uses shared memory without validating its contents, an attacker could manipulate the shared memory data and influence the execution of TEE applications.
Fig. 5 shows that variable buf is initialized by a shallow copy of the shared memory. Therefore, the contents of buf can be directly modified from outside TEE. When buf is passed into TEE_MemCompare function, the TEE application may not return the expected results. In order to solve this problem, shared memory should be deep copied to the in-TEE buffer, ensuring that the data used in TEE is isolated from external manipulations.
3. System Design
@@
In this section, we present the design of the proposed DITING to identify bad partitioning issues of TEE projects. With the help of CodeQL (Youn et al., 2023; GitHub, Inc., 2025), a code analysis engine developed by GitHub, we construct data-flows of parameters passed between the normal world and TEE based on Algorithms 2 to 4 in Section 3.3.1. Then, by analyzing the data-flows, DITING can detect insecure operations caused by bad partitioning. The design process involves 6 key steps shown in Fig. 6:
Figure 6. Overview workflow of DITING.
@@
Step 1: Parameter Type Identification. The initial step is identifying the parameters exchanged between the normal world and TEE. These parameters are classified into three types based on Section 2.1.
Step 2: Data-flow Generation. Data-flows of input/output and shared memory parameters will be generated. This will trace the path of each parameter and map its interactions with different components.
@@
Step 4: Rule Matching with Parameter Data-flows. If a parameter passes through rule-related components, we match its data-flow against the established rules in the preparation step to identify violations. Parameters that violate the rules will pinpoint bad partitioning issues, such as leakage of sensitive data to the normal world and misuse of shared memory leading to unexpected execution of in-TEE code.
@@
3.1. Rule Establishment
Given the discussion on bad partitioning in Section 2.2, we need to define a set of rules to detect these issues. These rules will focus on identifying incorrect parameter interactions with in-TEE sensitive functions and data that arise from bad partitioning.
Rule 1: Encrypting before Output. In this scenario, we mainly consider two critical functions — snprinf() and TEE_MemMove(), and assignment statements. The following code gives an implementation of secure in-TEE data output:
1char key[] = "123456";
2// buf is the ciphertext of key
3TEE_CipherUpdate(key, strlen(key), buf, &size);
4TEE_MemMove(params[0].memref.buffer, buf, size)
In line 4, key is encrypted by the TEE Internal API TEE_CipherUpdate() before copying to an output buffer via TEE_MemMove(dest, src, size), in which output params[0].memref.buffer is the destination buffer dest, and in-TEE data key is the source buffer src. Similarly, for the function snprintf(dest, size, format, args...), all arguments args must be encrypted before they are formatted to dest, which is a TEE output. Meanwhile, when an output value is the left value of an assignment statement, the corresponding right value should also be encrypted data. Rule 1 is shown as follow:
(1) $(e(buf)\Rightarrow w(o,buf))\land(e(v)\Rightarrow assgin(o,v))$
where $buf$ and $v$ are the contents of memory copy and assignment, $e()$ represents the encryption function, $w()$ and $assgin()$ will transfer $buf$ and $v$ to the output parameter $o$. The symbol $\Rightarrow$ defines that the left operation must be executed before the right operation.
Rule 2: Checking Value before in-TEE Use. According to the analysis in Section 2.2.2, even if the input value is checked in the normal world, the attacker still has the ability to tamper with the input and affect the execution of in-TEE code.
1char buf[] = "123456";
2unsigned int size = strlen(buf);
3if (params[0].memref.size < size) {
4 return TEE_ERROR_BAD_PARAMETERS;
@@ 3.1. Rule Establishment
6TEE_MemMove(params[0].memref.buffer, buf, size);
7...
8// a special case
9size = params[1].memref.size;
10char *str = TEE_Malloc(size, 0);
11TEE_MemMove(str, params[1].memref.buffer, size);
The above code provides the example of input value check. Before any input data is used, there should be a conditional statement containing that input. The usage of input values should be noted, including operations like indexing for array access, memory copying, and memory allocation. This leads us to define Rule 2:
(2) $isChecked(i)\Rightarrow array(i)\vee m(i)$
Here, $isChecked()$ denotes the input $i$ has been checked in a conditional statement; then, it can be used in array access $array()$ or memory operation functions $m()$, such as TEE_MemMove() and TEE_Malloc(). There is also a special case in lines 10 to 12. An memory space is allocated with the same size as the input buffer. Then, the input will be copied to the allocated memory. In this case, it is secure for an in-TEE dynamically allocated buffer based on the size of input buffer to accept input data, even if no conditional check is performed on the input. Therefore, the issue results need to filter this situation by the condition: $\neg(malloc(i)\Rightarrow m(i))$.
Rule 3: Using Deep Copy of Shared Memory. Since shared memory supports the feature of zero-copy, using a shallow copy of shared memory within TEE allows both the normal world and TEE to access the same physical memory. This creates a direct link between the trusted and untrusted environments, and breaks the isolation of TEE.
1unsigned int size = params[1].memref.size;
2char *buf = TEE_Malloc(size);
3TEE_MemMove(buf, params[1].memref.buffer, size);
4...
5if (!strcmp(buf, "from_CA_to_TA")) {
6 ...
7}
As shown in the above code, the first step after transferring a shared memory buffer into TEE should be creating a deep copy of the shared memory. This ensures that TEE operates on an independent copy of the data, maintaining its isolation and preventing any unauthorized access or influence from the normal world. In Rule 3, an in-TEE buffer $buf$ needs to be initialized via function $deepCopy()$ with shared memory $shm$. Then, any memory operations $m()$ could be performed on $buf$:
(3) $deepCopy(buf,shm)\Rightarrow m(buf)$
3.2. Parameter Type Identification
Since DITING requires constructing and analyzing data-flows for different types of parameters, it needs to first identify the type of each parameter transferred between the normal world and TEE. For example, in TrustZone-based OP-TEE, there are 10 parameter names, and we classify them into 3 types, which are listed in Table 3.
Table 3. Types of parameters in TrustZone-based OP-TEE.
Name Type
TEEC_VALUE_OUTPUT Output
TEEC_MEMREF_OUTPUT
TEEC_VALUE_INPUT Input
TEEC_MEMREF_INPUT
TEEC_VALUE_INOUT Output & Input
TEEC_MEMREF_INOUT
TEEC_MEMREF_WHOLE Shared Memory
TEEC_MEMREF_PARTIAL_INPUT
TEEC_MEMREF_PARTIAL_OUTPUT
TEEC_MEMREF_PARTIAL_INOUT
Input: $F$, all function TEEC_InvokeCommand() calls;
          $S$, all definition statements of parameter type
Output: $T$, types of parameters passed into each function call $f$ of $F$
1 foreach $f\in F$ do
       // $T_{f}$ means the f-th element in $T$, and $S_{0}$ is the first element of $S$
2       $T_{f}\leftarrow S_{0}$;
3       foreach $s\in S$ do
4             if $s$.isBefore($f$) $\land$ $T_{f}$.isBefore($s$) then
5                   $T_{f}\leftarrow s$;
6
7
@@ 5                   $T_{f}\leftarrow s$;
Algorithm 1 Determine the parameter types passed into OP-TEE.
1op.paramTypes = TEEC_PARAM_TYPES(...);
2res = TEEC_InvokeCommand(&session, TA_CMD, &op, &err_origin);
The above code shows that the parameter types are defined by the line 1 statement before the TEE invocation function TEEC_InvokeCommand() in line 2. Since the normal-world code may contain multiple pairs of parameter definition and TEE invocation, for a TEEC_InvokeCommand() call, we need to locate the parameter type definition statement that precedes it and is closest to it. Algorithm 1 gives the determination of the parameter types passed into one invocation. The algorithm iterates through all parameter definitions and TEE invocation statements, identifying the corresponding parameter definition statement for the current TEE invocation according to the condition in line 4, which means there are no other parameter definition statements between a parameter definition and the TEE invocation. Then, we can generate the data-flows of these parameters by code analysis tools, such as CodeQL.
3.3. Detection of Bad Partitioning
3.3.1. Data-flow Generation
In this step, we first build the data-flow of each parameter passed between the normal world and TEE.
Table 4. The sources and sinks of data-flow generation for output parameters.
Sources (1) in-TEE sensitive data
Sinks (1) output buffer, which is the destination buffer for copying the source: $w(o,source)$
(2) output value, which is assigned with the source: $o=source$
(3) all function calls, in which the source is passed: $functionCall(source)$
Input: $F$, all function calls;
          $S$, all assignment statements
Output: $D$, sensitive data list;
            $ODF$, data-flows for outputs
1 $D\leftarrow\emptyset$;
2 $ODF\leftarrow\emptyset$;
3 foreach $f\in F$ do
       // $f$ has the dest argument, which is an output
4       if $f$.hasDestArgument() is output then
             // get the data copied to output
5             $D\leftarrow D\cup\{f.\text{{{geSrcArgument(}}{\emph{}}{{)}}}\}$( ) };
6
7
8foreach $s\in S$ do
       // get the assigned object
9       if $s$.getLeftValue() is output then
             // get the data assigned to output
10             $D\leftarrow D\cup\{s.\text{{{getRightValue(}}{\emph{}}{{)}}}\}$( ) };
@@ 2 $ODF\leftarrow\emptyset$;
13foreach $d\in D$ do
14       foreach $f\in F$ do
             // check whether the data passed in $f$
15             if $f$.hasArgument($d$) is True then
16                   $ODF_{d}\leftarrow ODF_{d}\cup\{f\}$;
17
18
19
Algorithm 2 Data-flow generation for output parameters.
Table 5. The sources and sinks of data-flow generation for input parameters.
Sources (1) input value
(2) input buffer
Sinks (1) conditional statement with the source: $if(source)$
(2) index of array access, which is the source or expression with the source: $array(source)\vee array(expr(source))$
(3) memory operation function, in which the source passed: $m(source)$
Input: $I$, all input parameters;
          $C$, all conditional statements;
          $A$, all array access statements;
          $F$, all function calls;
Output: $IDF$, data-flows for inputs
1 $IDF\leftarrow\emptyset$;
2 foreach $i\in I$ do
3       foreach $c\in C$ do
             // $i$ is in the controlling exprseeion $e$ of $c$
4             $e\leftarrow c$.getControllingExpr();
5             if $e$.hasOperand($i$) is True then
6                   $IDF_{i}\leftarrow IDF_{i}\cup\{c\}$;
7
8
9      foreach $a\in A$ do
             // $i$ is an index of $a$
10             if $a$.getIndex() is $i$ then
11                   $IDF_{i}\leftarrow IDF_{i}\cup\{a\}$;
12
            // check whether $i$ is in an expression
13             if isInExpr($i$) is True then
                   // get the expression where the $i$ is an operand
14                   $e\leftarrow i$.getExpr();
15                   if $a$.getIndex() is $e$ then
16                         $IDF_{i}\leftarrow IDF_{i}\cup\{a\}$;
17
18
19
20      foreach $f\in F$ do
21             if $f$.hasArgument($i$) is True then
22                   $IDF_{i}\leftarrow IDF_{i}\cup\{f\}$;
23
24
25
Algorithm 3 Data-flow generation for input parameters.
For output parameters, the sources and sinks of their data-flow generation are listed in Table 4. Sensitive data in TEE, such as cryptographic keys and authentication credentials, will be the sources. The sinks consist of output parameters that serve as the targets of source copies and assignment operations. In addition, we also need to obtain what operations the sensitive data in TEE has undergone, that is, which functions they have been called into. Algorithm 2 describes the construction of data-flows for outputs, it has 2 stages:
  1. a)
    Lines 3 to 8 are the procedure of obtaining all in-TEE data that will be copied or assigned to the output parameters.
  2. b)
    Lines 9 to 12 can get all functions that have processed the above data.
Then, we can trace the flow paths of in-TEE data to the outputs $ODF$, and $ODF_{d}$ means the path of $d$.
For input parameters, as illustrated in Table 5, we only need to focus on the statements that they pass through, such as conditions, array access, memory operation function calls, etc. Additionally, the second sink needs to consider two types: a[i] and a[i-1], which are i and the expression composed of i as an array index, respectively. There are three loop blocks at lines 3 to 16 in Algorithm 3. Each of them corresponds to three kinds of sinks in Table 5. Finally, we get $IDF$ which represents the data-flow of each input during execution of in-TEE code, and $IDF_{i}$ denotes data-flow path of $i$.
In the end, the sinks of shared memory data-flows are the buffer initialization statements based on a shallow copy of shared memory, and function calls where they directly treat the shared memory buffer as the argument. According to Table 6, we design Algorithm 4 to generate the data-flows of shared memory parameters. Lines 3 to 5 are responsible for retrieving the status of the first sink, lines 6 to 8 correspond to the second, and $MDF_{m}$ is data-flow of $m$.
Table 6. The sources and sinks of data-flow generation for shared memory.
Sources (1) shared memory buffer
Sinks (1) shallow copy statement with the source: $buf=shm$
(2) all function calls, where directly use the source: $f(shm)$
Input: $M$, all shared memory parameters;
          $S$, all assignment statements;
          $F$, all function calls;
Output: $MDF$, data-flows for shared memory
1 $MDF\leftarrow\emptyset$;
2 foreach $m\in M$ do
3       foreach $s\in S$ do
4             if $s$.getRightValue() is $m$ then
5                   $MDF_{m}\leftarrow MDF_{m}\cup\{s\}$;
6
7
8      foreach $f\in F$ do
9             if $f$.hasArgument($m$) is True then
10                   $MDF_{m}\leftarrow MDF_{m}\cup\{f\}$;
11
12
13
Algorithm 4 Data-flow generation for shared memory.
3.3.2. Rule Matching and Issue Reporting
Once the data-flow of each parameter is constructed, it can be matched against the rules defined in Section 3.1 to identify bad partitioning issues in TEE projects.
Bad partitioning ①. Algorithm 5 describes a way of detecting unencrypted data that will be written to the output. First, for each data-flow path in $ODF$, we initialize a flag $isEncrypted$ to False. This flag tracks whether data in the current path has been encrypted. Then, we iterate through each node $n$ in the current path. If $n$ represents a function call, we need to extract its name and check whether the name contains substrings like enc or cipher. If such a substring is present, set $isEncrypted$ to True. This ensures encryption functions within a data-flow are identified to determine the data has been protected.
For other function nodes, if they have a destination buffer argument that is an output, we should check whether $isEncrypted$ is False. If so, we can record this node in the report $OR$. For $n$, which is an assignment statement, if the left-value of this assignment is an output and $isEncrypted$ of this path is still False while iterating to the node, this node also needs to be put into $OR$. After processing all paths in $ODF$, Algorithm 5 returns the report $OR$, which contains all nodes where unencrypted data is copied to the output.
Input: $ODF$, data-flows for outputs
Output: $OR$, report of unencrypted data output
1 foreach $path\in ODF$ do
2       $isEncrypted\leftarrow False$;
       // $n$ is a node in the $path$
3       foreach $n\in path$ do
4             if isFunction($n$) is True then
                   // get the function name of $n$
5                   if (``$enc"\vee``cipher")$ in $n$.name() then
6                         $isEncrypted\leftarrow True$;
7
8                  if $n$.hasDestArgument() is output then
9                         if $isEncrypted$ is False then
10                               $OR\leftarrow OR\cup\{n\}$;
11
12
13
14            if isAssignment($n$) is True then
15                   if $n$.getLeftValue() is output then
16                         if $isEncrypted$ is False then
17                               $OR\leftarrow OR\cup\{n\}$;
18
19
20
21
22
Algorithm 5 Detection of unencrypted data output.
Bad partitioning ②. Algorithm 6 focuses on detecting potential weaknesses in input validation by checking whether inputs pass through conditional logic before being processed further. It assumes that conditional statements are indicators of input validation and marks nodes of $IDF$ as vulnerable if they precede any such checks.
Similar to Algorithm 5, a flag $isChecked$ is also initialized to False and we retrieve each node in paths of $IDF$. The flag indicates whether input validation has been performed along the path. If the node represents a conditional statement (e.g., an if or switch block), we can set $isChecked$ to True. If $isChecked$ remains False for a node $n$, it suggests that no validation has been performed for this input up to that point. The node $n$ will be added to the report $IR$. In addition, we exclude the special case in Rule 2 through lines 9 to 14.
Input: $IDF$, data-flows for inputs
Output: $IR$, report of input validation weaknesses
1 foreach $path\in IDF$ do
2       $isChecked\leftarrow False$;
3       $isSecure\leftarrow False$;
4       foreach $n\in path$ do
5             if isConditional($n$) is True then
6                   $isChecked\leftarrow True$;
7
8            if $isChecked$ is False then
9                   $IR\leftarrow IR\cup\{n\}$;
10
            // filter the special case
11             if isFunction($n$) is True then
12                   if (``$malloc")$ in $n$.name() then
13                         $isSecure\leftarrow True$;
14
15                  if (``$memove")$ in $n$.name() then
16                         if $isSecure$ is True then
17                               $IR\leftarrow IR-\{n\}$;
18
19
20
21
22
Algorithm 6 Detection of input validation weaknesses.
Bad partitioning ③. Algorithm 7 identifies direct access to shared memory without secure-guards, which can lead to in-TEE data corruption and leakage, or other issues of data integrity and confidentiality. Rule 3 ensures that shared memory is only used after appropriate handling, such as deep copying, to avoid effects from the normal world.
If a node $n$ represents a function call, we will determine whether it performs a deep copy operation where the shared memory $shm$ is the source buffer. If it is true, we could continue to check the next node. For all other cases where $shm$ is directly accessed, then, the node $n$ is tagged on the bad partitioning of directly using shared memory.
Input: $MDF$, data-flows for shared memory
Output: $MR$, report of direct usage of shared memory
1 foreach $path\in MDF$ do
2       foreach $n\in path$ do
3             if isFunction($n$) is True then
                   // $n$ is a deep copy where shared memory is source buffer
4                   if $n$.hasSrcArgument() is $shm$ then
5                         continue;
6
7
8            $MR\leftarrow MR\cup\{n\}$;
9
10
Algorithm 7 Detection of direct usage of shared memory.
Finally, three kinds of bad partitioning within TEE projects are reported by DITING. Moreover, DITING can provide strong scalability; by simply adding new rules to the system, it can be extended to detect a broader range of potential problems. Thence, DITING could adapt to evolving requirements and uncover additional vulnerabilities arisen from bad partitioning, making it valuable for ensuring robust and secure implementations of TEE projects.
4. Evaluations

### v2

@@ II-C3 Direct Usage of Shared Memory
Shared memory, as a zero-copy memory block, provides the same physical memory space accessible to both the secure and non-secure partitions. Since this memory is maintained by the normal-world code, directly using shared memory can introduce significant vulnerabilities. Fig. 5 shows that variable buf is initialized by a shallow copy of the shared memory. Therefore, the contents of buf can be directly modified from outside TEE. When buf is passed into TEE_MemCompare function, the TEE application may not return the expected results. Therefore, we consider the code that directly uses shared memory to be an insecure component.
Figure 6: Overview workflow of DITING.
III System Design
@@
In this section, we present the design of the proposed DITING  [^5] https://github.com/CharlieMCY/PartitioningE-in-TEE to identify bad partitioning issues of TEE projects. With the help of CodeQL [28, 29], a code analysis engine developed by GitHub, we construct data flows of parameters passed between the normal world and TEE. Then, by analyzing the data flows, DITING can detect insecure operations caused by bad partitioning. The design process involves 6 key steps shown in Fig. 6:
@@
Step 1: Parameter Type Identification. The initial step is identifying the parameters exchanged between the normal world and TEE. These parameters are classified into three types based on Section II-A.
Step 2: Data flow Generation. Data flows of input/output and shared memory parameters will be generated. This will trace the path of each parameter and map its interactions with different components.
@@
Step 4: Rule Matching with Parameter Data flows. If a parameter passes through rule-related components, we match its data flow against the established rules in the preparation step to identify violations. Parameters that violate the rules will pinpoint bad partitioning issues, such as leakage of sensitive data to the normal world and misuse of shared memory leading to unexpected execution of in-TEE code.
@@
III-A Rule Establishment
TABLE II: Defined notations of parameters, variables, and operations in the TEE code.
$o$ : The output parameter
$i$ : The input parameter
$m$ : The shared memory parameter
$x,y,z$ : The data in the TEE code
$y:=\texttt{ENC}(x)$ : $y$ is a return value or buffer of the encryption-related cryptographic API, where $x$ is passed into. It can also be said that $x$ is encrypted into cipher text $y$
$\texttt{COPY}(d,s)$ : Copying memory from the source buffer $s$ to destination buffer $d$
$\texttt{IF}(i)$ : $i$ exists in the expression of a conditional statement
$\texttt{ARRAY}(i)$ : An $i$-related array access
$x:=\texttt{SHALLOW}(m)$ : $x$ is the shallow copy of the shared memory parameter $m$
$x:=\texttt{MALLOC}(i)$ : An in-TEE buffer $x$ is allocated according to the input parameters $i$
$y:=\texttt{OP}(x)$ : Other operations besides the above, including arithmetic, function passing…
Given the discussion on bad partitioning in Section II-C, we need to define a set of rules to detect these issues. These rules will focus on identifying insecure parameter interactions with in-TEE sensitive functions and data, which commonly arise from bad partitioning. As shown in Table II, we use lowercase letters ($o$, $i$, $m$, $x$…) to denote parameters and data in the TEE code. Meanwhile, we define some commonly used operations in the TEE code, such as data encryption functions $y:=\texttt{ENC}(x)$, conditional statements $\texttt{IF}(i)$, access to array $\texttt{ARRAY}(i)$, shallow copying $x:=\texttt{SHALLOW}(m)$, etc. In addition, it should be noted that $\texttt{COPY}(d,s)$ is a deep copy function, which has two situations: output to the outside of TEE $\texttt{COPY}(o,x)$ and copying data into TEE $\texttt{COPY}(x,i)$.
Based on these notations, we provide the rules in Fig. 7 and use the predicates in Table III for generating the data flows and detecting the issues. Firstly, we give the propagation rules to represent the transfer path of in-TEE data or parameters. $x\rightsquigarrow y$ denotes direct propagation, indicating that $y$ is directly derived from $x$, for example, $x$ is passed to $y$ through a function call or assignment statement. Then, $x\rightsquigarrow^{*}z$ denotes a transitive propagation, meaning that $x$ has undergone multiple propagations to $z$.
[…]
[Transitive Propagation] x ↝y
y ↝z x ↝^* z
[…]
[EncStmt] y := ENC(x) SC(x)
[…]
[Sink] y ↝^* z
COPY(o,z) Sink(o)
[…]
[Violation] Sink(o)
SC(x) BP
Rule 2 for Input Validation Weaknesses {mathpar}
[…]
[Source] i Source(i)
[…]
[IfStmt] IF(i) SC(i)
[…]
[Sink] i ↝^* i’
CS∈{ARRAY(i’),COPY(x,i’)} Sink(CS)
[…]
[Violation] Sink(CS)
SC(i) BP
[…]
[ShallowCopyStmt] x := SHALLOW(m) ISC(x)
ISC(m)
[…]
[Sink] m ↝^* y
x ↝^* y Sink(y)
[…]
[Violation] Sink(y)
ISC(m) ∨∃ISC(x) BP
Figure 7: Rules for detecting bad partitioning (BP) issues by checking the violation usage of TEE parameters. The violations include two types: (1) sensitive statements or parameter usage that lack security components, (2) direct usage of insecure components.
TABLE III: Predicates used in the static analysis. * denotes a node in the data flow, and the predicate presents the state of a node based on the rules.
Predicates Description
$\texttt{Source}(*)$ is a source node of the data flow
$\texttt{Sink}(*)$ is a sink node of the data flow
$\texttt{SC}(*)$ is a secure component
$\texttt{ISC}(*)$ is a insecure component
CS A critical statement
BP A bad partitioning issue
In Rule 1, we assume that $x$ is an in-TEE data that is eventually passed to an output parameter. If $x$ is not passed into an encryption API (e.g., TEE_CipherUpdate()), it is treated as unencrypted (i.e., plain text). Consequently, any output parameters that receive $x$ via memory copy operations (e.g., $\texttt{COPY}(o,x)$) are identified as bad partitioning instances.
[…]
COPY(x,i’)
∃x:=MALLOC(i’) ¬BP
Rule 3 captures a scenario where shallow copying shared memory allows both the TEE and the normal world to access the same physical memory region. This violates the isolation guarantee of TEE and is therefore identified as a bad partitioning instance under our threat model.
To facilitate the detection of bad partitioning issues, we further refine the preceding rules by introducing a fine-grained classification of TEE operations. Specifically, memory copy and array access operations are defined as critical statements (CS), while encryption functions and conditional checks are classified as secure components (SC). Meanwhile, the TEE code that directly accesses normal-world memory (i.e., shared memory) is an insecure component (ISC). Based on this classification:
  • Violations of Rules 1 and 2 can be detected by identifying critical statements that lack an associated secure component;
  • Violations of Rule 3, which involve the direct use of shared memory or its shallow copies, can be identified by detecting the presence of an insecure component.
III-B Data flow Generation
In this step, we construct the data flow of each parameter passed between the normal world and TEE according to the source and sink shown in Fig. 7. For the data flow of output parameters, the source is the in-TEE data, and the sink is an output parameter, which is the destination buffer for a memory copy function or the left value of an assignment statement. For input parameters, the source is the input value provided by the normal world, and the sink includes input-related operations such as array accesses and memory copies. For shared memory, the data flow encompasses all operations that directly access shared memory or its shallow copies within the TEE code.
III-C Detection of Bad Partitioning
Input: $DF$, the data flow of TEE parameter
Output: $R$, the report of bad partitioning issues
1 $isBP\leftarrow True$;
2 foreach $node\in DF$ do
3       if isSecureComponent($node$) then
4             $isBP\leftarrow False$;
@@
6      if isInsecureComponent($node$) then
7             $R\leftarrow R\cup\{node\}$;
@@
9      if isCriticalStmt($node$) and $isBP$ then
10             $R\leftarrow R\cup\{node\}$;
@@
Algorithm 1 Detection of bad partitioning issues based on the data flow.
Based on the rules in Section III-A and data flow in Section III-B, we can detect bad partition issues by traversing the data flow graph and checking rule violations at each relevant node. As shown in Algorithm 1, each $node$ in the data flow $DF$ presents an operation statement, and all nodes are ordered according to program execution. A node is flagged as a bad partitioning instance if the node is a critical statement that is not preceded by a corresponding secure component as required by the rules. Additionally, any insecure component (e.g., direct use of shared memory) is immediately reported as a violation. Since the data flow may involve different types of parameters (input, output, or shared memory), each detected issue is recorded in the relevant report $R$ according to its associated parameter type.
IV Evaluations